AVX Home Entertainment, Inc., recently began a “no-hassles” return policy. A sample of 540 customers who recently returned items showed 360 thought the policy was fair, 150 thought it took too long to complete the transaction, and the rest had no opinion. On the basis of this information, make an inference about customer reaction to the new policy. (Round your answers to 1 decimal place.)

fair percentage:_______ too long percentage:_______ no opinion percentage:________
Answer #1

Given sample n=540

Fair percentage= no. Of customers thought fair / n

=360/540

=0.6667=66.7%

Too long percentage = 150/540

=0.2778=27.8%

No opinion Customer's are =540-360-150 = 30

No opinion percentage=30/540

=0.0556=5.6%